OpenStudy (anonymous):

Help plzz FANS & MEDALS!! What is the best way to classify each equation? A. identity B. contradiction C. neither

OpenStudy (anonymous):

1.5x-20=2(3x-10) 2.4x-16-2x=2(x-8) 3.6x-15=3(2x-6) 4.5x-3(x-4)=2(x-6)